OpenStudy (anonymous):

What is the simplified form of 1 over x squared all over 3 over x-cubed ? x over 3 1 over 3x <---- MY ANSWER 3x 3 over x @satellite73 Can u plz check my answer?

OpenStudy (anonymous):

\[\frac{1}{x^2}\times \frac{x^3}{3}\]m

OpenStudy (anonymous):

final answer will be \[\frac{x}{3}\]

OpenStudy (anonymous):

by cancelling or as you might say "x over 3"

OpenStudy (anonymous):

oh so u flip division into multiplication? @satellite73

OpenStudy (anonymous):

yes, just like with numbers

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

\[\huge \frac{\frac{a}{b}}{\frac{c}{d}}=\frac{ad}{bc}\]

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Thank youu!!

OpenStudy (anonymous):

yw